# PerfGuard: A Performance-Aware Agent for Visual Content Generation

### v1

@@ Abstract
footnotetext: * Corresponding author: like1990@bupt.edu.cn
@@ 2 Relative Works
Recent advances in visual content generation have significantly improved controllability and semantic alignment. Models like FLUX (Labs, 2024), Stable Diffusion3 (Esser et al., 2024), and DALL·E3 (Betker et al., 2023) generate images from textual prompts, while ControlNet (Zhang et al., 2023), T2I-Adapter (Mou et al., 2024), and InstanceDiffusion (Wang et al., 2024a) incorporate multimodal signals to better match user intent. To support fine-grained control, LayoutGPT (Feng et al., 2023), RPG (Yang et al., 2024b), GoT (Fang et al., 2025a), and T2I-R1 (Jiang et al., 2025) leverage LLMs to decompose prompts into region-specific semantics. Systems like CompAgent (Wang et al., 2024c) and GenArtist (Wang et al., 2024b) coordinate generation and editing tools, while MCCD (Li et al., 2025) and T2I-Copilot (Chen et al., 2025) improve performance through model cooperation. CLOVA (Gao et al., 2024a) improves the success rate of visual tasks, including face swapping, by enhancing the tool’s prompt pool through the introduction of self-reflection and prompt tuning. However, most approaches assume reliable tool execution and overlook how performance boundaries affect planning accuracy. PerfGuard addresses this gap by explicitly modeling tool capabilities and execution feedback.
@@ 5 Experiments
We conducted both qualitative and quantitative comparisons of PerfGuard against various image generation and editing models. The evaluation spans three benchmarks covering different task types: basic image generation (T2I-CompBench (Huang et al., 2023)), advanced image generation (OneIG-Bench (Chang et al., 2025)), and complex image editing (Complex-Edit (Yang et al., 2025b)). Detailed experimental settings, descriptions of baseline methods, agent prompts and instructions, as well as additional results and visualizations, are provided in the supplementary material A.10..
@@ 5.1 Qualitative Results and Analysis
We compared the proposed PerfGuard with several existing methods on text-to-image generation and image editing tasks. The visualization results reveal three key observations: i) In text-to-image generation, traditional diffusion models struggle with complex prompts involving multiple entities and detailed attributes. Their limited language understanding leads to poor semantic alignment. For example, FLUX (Labs, 2024) and SD3 (Esser et al., 2024) fail to generate a cat in a spacesuit. CoT-based methods like T2I-R1 (Jiang et al., 2025) and GoT (Fang et al., 2025a) incorporate LLMs, but due to reliance on a single-generation tool, they still miss key elements or actions, such as GoT omitting the fishing pose and several specified objects. Agent-based methods show improvement in semantic parsing and tool orchestration. However, GenArtist (Wang et al., 2024b) lacks a performance-aware tool selection strategy, resulting in planning errors and missing elements. T2I-Copilot (Chen et al., 2025) performs better through multi-module semantic decomposition, but its limited tool diversity still leads to omissions, such as spiral galaxies and green glasses. ii) In multi-round editing tasks, traditional methods like ICEdit (Zhang et al., 2025c) and AnySD (Yu et al., 2025) deliver the weakest results. GenArtist, despite using multiple tools, suffers from poor capability matching, leading to suboptimal edits. Step1X_Edit (Liu et al., 2025) benefits from LLM-enhanced understanding of long instructions, but without intelligent planning and execution, it fails to capture key details—for example, the kite does not transform into a dragon. iii) Across both generation and editing tasks, PerfGuard consistently achieves the most accurate and visually aligned outputs. This demonstrates that its performance-guided tool selection enhances single-step execution accuracy and improves overall task planning.
@@ 5.2 Quantitative Results and Analysis
T2I-Copilot (Chen et al., 2025) Agent 0.821 0.679 0.318 0.386
@@ 5.4 Efficiency Comparison
Token Consumption Comparison   To demonstrate the efficiency of our method in tool selection, we expand the problem into large-scale tool management within future agent communities. Specifically, we simulate a large tool library using GPT-4o (Fang et al., 2025b), where the number of tools ranges from 10 to 200, generating tool information with textual descriptions and multi-dimensional ratings. The specific details are provided in the Supplementary Material A.9. We use the “complex_vel” subset from T2I-CompBench for the task and compare PerfGuard’s performance-driven tool selection with traditional text-based methods, with a maximum token output of 8192. We compare total token consumption (input and output) between the two methods. Fig. 8 shows: 1) The traditional method consumes more tokens, as it struggles to define tool capabilities, resulting in catastrophic growth in token consumption as the number of tools increases, without addressing selection correctness. 2) PerfGuard, by focusing on task-specific dimensions, is unaffected by the number of tools. 3) As dimensions increase (from d=4 to d=16), token consumption for PerfGuard mainly increases slowly in the input prompts. This demonstrates PerfGuard’s superior efficiency in tool management and selection for future agents.
6 Limitations and Challenges
@@ 7 Conclusion
We are committed to ensuring the reproducibility of our work. To facilitate this, we have provided comprehensive implementation details, experimental settings, and evaluation protocols in the main paper and appendix. All datasets used in our experiments are publicly available, and we include detailed data preprocessing steps in the supplementary materials. For our proposed framework and algorithms, we have submitted an anonymous source code repository as part of the supplementary materials, which includes scripts for training, evaluation, and visualization. The repository is available at https://github.com/FelixChan9527/PerfGuard_agent. We also provide ablation studies and hyperparameter configurations to support reproducibility. We encourage readers and reviewers to refer to the appendix and supplementary files for further details.
@@ A.2 EXPERIMENTAL SETUP
Competitors  i) For the image generation task, we systematically compared three categories of methods: diffusion model-based approaches (e.g., FLUX (Labs, 2024), SD3 (Esser et al., 2024)), Chain-of-Thought (CoT)-based approaches (e.g., GoT (Fang et al., 2025a), T2I-R1 (Jiang et al., 2025)), and agent-based approaches (e.g., GenArtist (Wang et al., 2024b), T2I-Copilot (Chen et al., 2025)). By contrasting these strategies, we aim to analyze how different visual reasoning mechanisms impact the semantic accuracy of generated images. ii) For the image editing task, we evaluated not only pure diffusion-based methods (e.g., ICEdit (Zhang et al., 2025c), AnySD (Yu et al., 2025)) but also Step1X_Edit (Liu et al., 2025), which integrates large language model (LLM) techniques. To ensure a fair comparison, we additionally included general-purpose models capable of both image generation and editing (e.g., GenArtist and OmniGen (Xiao et al., 2025)).
@@ A.5 Ablation study with different LLMs
A.6 Design and Specification of Performance Boundaries in PerfGuar
@@
A.7 Definition of $\mathcal{R}_{\text{actual}}$ in Adaptive Preference Updating
@@ A.8 Prompt Engineering
A.9 Configurations Related to the Tool Selection Simulation Experiment in Sec. 8
@@
A.10 More visualization results

### v2

@@ Abstract
footnotetext: * Chao Zhang and Ke Li are co-corresponding authors: ariczhang2009@gmail.com, like1990@bupt.edu.cn
@@ 2 Relative Works
Recent advances in visual content generation have significantly improved controllability and semantic alignment. Models like FLUX (Labs, 2024), Stable Diffusion3 (Esser et al., 2024), and DALL·E3 (Betker et al., 2023) generate images from textual prompts, while ControlNet (Zhang et al., 2023), T2I-Adapter (Mou et al., 2024), InstanceDiffusion (Wang et al., 2024a) and VersaGen (Chen et al., 2025b) incorporate multimodal signals to better match user intent. To support fine-grained control, LayoutGPT (Feng et al., 2023), RPG (Yang et al., 2024b), GoT (Fang et al., 2025a), and T2I-R1 (Jiang et al., 2025) leverage LLMs to decompose prompts into region-specific semantics. Systems like CompAgent (Wang et al., 2024c) and GenArtist (Wang et al., 2024b) coordinate generation and editing tools, while MCCD (Li et al., 2025) and T2I-Copilot (Chen et al., 2025a) improve performance through model cooperation. CLOVA (Gao et al., 2024a) improves the success rate of visual tasks, including face swapping, by enhancing the tool’s prompt pool through the introduction of self-reflection and prompt tuning. However, most approaches assume reliable tool execution and overlook how performance boundaries affect planning accuracy. PerfGuard addresses this gap by explicitly modeling tool capabilities and execution feedback.
@@ 5 Experiments
We conducted both qualitative and quantitative comparisons of PerfGuard against various image generation and editing models. The evaluation spans three benchmarks covering different task types: basic image generation (T2I-CompBench (Huang et al., 2023)), advanced image generation (OneIG-Bench (Chang et al., 2025)), and complex image editing (Complex-Edit (Yang et al., 2025b)). Detailed experimental settings, descriptions of baseline methods, agent prompts and instructions, as well as additional results and visualizations, are provided in the supplementary material A.11..
@@ 5.1 Qualitative Results and Analysis
We compared the proposed PerfGuard with several existing methods on text-to-image generation and image editing tasks. The visualization results reveal three key observations: i) In text-to-image generation, traditional diffusion models struggle with complex prompts involving multiple entities and detailed attributes. Their limited language understanding leads to poor semantic alignment. For example, FLUX (Labs, 2024) and SD3 (Esser et al., 2024) fail to generate a cat in a spacesuit. CoT-based methods like T2I-R1 (Jiang et al., 2025) and GoT (Fang et al., 2025a) incorporate LLMs, but due to reliance on a single-generation tool, they still miss key elements or actions, such as GoT omitting the fishing pose and several specified objects. Agent-based methods show improvement in semantic parsing and tool orchestration. However, GenArtist (Wang et al., 2024b) lacks a performance-aware tool selection strategy, resulting in planning errors and missing elements. T2I-Copilot (Chen et al., 2025a) performs better through multi-module semantic decomposition, but its limited tool diversity still leads to omissions, such as spiral galaxies and green glasses. ii) In multi-round editing tasks, traditional methods like ICEdit (Zhang et al., 2025c) and AnySD (Yu et al., 2025) deliver the weakest results. GenArtist, despite using multiple tools, suffers from poor capability matching, leading to suboptimal edits. Step1X_Edit (Liu et al., 2025) benefits from LLM-enhanced understanding of long instructions, but without intelligent planning and execution, it fails to capture key details—for example, the kite does not transform into a dragon. iii) Across both generation and editing tasks, PerfGuard consistently achieves the most accurate and visually aligned outputs. This demonstrates that its performance-guided tool selection enhances single-step execution accuracy and improves overall task planning.
@@ 5.2 Quantitative Results and Analysis
T2I-Copilot (Chen et al., 2025a) Agent 0.821 0.679 0.318 0.386
@@ 5.4 Efficiency Comparison
Token Consumption Comparison   To demonstrate the efficiency of our method in tool selection, we expand the problem into large-scale tool management within future agent communities. Specifically, we simulate a large tool library using GPT-4o (Fang et al., 2025b), where the number of tools ranges from 10 to 200, generating tool information with textual descriptions and multi-dimensional ratings. The specific details are provided in the Supplementary Material A.10. We use the “complex_vel” subset from T2I-CompBench for the task and compare PerfGuard’s performance-driven tool selection with traditional text-based methods, with a maximum token output of 8192. We compare total token consumption (input and output) between the two methods. Fig. 8 shows: 1) The traditional method consumes more tokens, as it struggles to define tool capabilities, resulting in catastrophic growth in token consumption as the number of tools increases, without addressing selection correctness. 2) PerfGuard, by focusing on task-specific dimensions, is unaffected by the number of tools. 3) As dimensions increase (from d=4 to d=16), token consumption for PerfGuard mainly increases slowly in the input prompts. This demonstrates PerfGuard’s superior efficiency in tool management and selection for future agents.
@@ 6 Conclusion
Acknowledgments
This work was supported in part by the Postdoctoral Fellowship Program of the China Postdoctoral Science Foundation (CPSF) under Grant No. GZC20251088.
@@ Acknowledgments
We are committed to ensuring the reproducibility of our work. To facilitate this, we have provided comprehensive implementation details, experimental settings, and evaluation protocols in the main paper and appendix. All datasets used in our experiments are publicly available, and we include detailed data preprocessing steps in the supplementary materials. For our proposed framework and algorithms, we have submitted an anonymous source code repository as part of the supplementary materials, which includes scripts for training, evaluation, and visualization. The repository is available at https://github.com/FelixChan9527/PerfGuard. We also provide ablation studies and hyperparameter configurations to support reproducibility. We encourage readers and reviewers to refer to the appendix and supplementary files for further details.
@@ A.2 EXPERIMENTAL SETUP
Competitors  i) For the image generation task, we systematically compared three categories of methods: diffusion model-based approaches (e.g., FLUX (Labs, 2024), SD3 (Esser et al., 2024)), Chain-of-Thought (CoT)-based approaches (e.g., GoT (Fang et al., 2025a), T2I-R1 (Jiang et al., 2025)), and agent-based approaches (e.g., GenArtist (Wang et al., 2024b), T2I-Copilot (Chen et al., 2025a)). By contrasting these strategies, we aim to analyze how different visual reasoning mechanisms impact the semantic accuracy of generated images. ii) For the image editing task, we evaluated not only pure diffusion-based methods (e.g., ICEdit (Zhang et al., 2025c), AnySD (Yu et al., 2025)) but also Step1X_Edit (Liu et al., 2025), which integrates large language model (LLM) techniques. To ensure a fair comparison, we additionally included general-purpose models capable of both image generation and editing (e.g., GenArtist and OmniGen (Xiao et al., 2025)).
@@ A.5 Ablation study with different LLMs
A.6 Limitations and Challenges
A.7 Design and Specification of Performance Boundaries in PerfGuar
@@
A.8 Definition of $\mathcal{R}_{\text{actual}}$ in Adaptive Preference Updating
@@ A.9 Prompt Engineering
A.10 Configurations Related to the Tool Selection Simulation Experiment in Sec. 8
@@
A.11 More visualization results